C/D or C over D + 2
C/D+2 = ???
Answers are C+2D over D or C+2D

Also what are the possible values of x such that 3(xsquared) - 2x =0
three x squared minus two x =0
-2/3 or 0 or 2/3 or 0 and 2/3
finally,
sqrt 3 + sqrt 27 = ?????


Answer by stanbon(75887)

(C/D)+2 = ???
Answers are C+2D over D or C+2D
least common denominator is "D":
= (C/D) + (2D/D)
= (c+2D)/D
---------------------------------

Also what are the possible values of x such that 3(x^2) - 2x =0
Factor to get:
x(3x-2) = 0
Since the product is zero, one of the factors must be zero:
x = 0 or 3x-2 = 0
x = 0 or x = 2/3
------------------------
finally,
sqrt 3 + sqrt 27 = ?????
= sqrt(3) + sqrt(9*3)
= sqrt(3) + sqrt(9)*sqrt(3)
= sqrt(3) + 3*sqrt(3)
= 4sqrt(3)
===============
Cheers,
Stan H.